What is the probability that two draws have different colors?

What is the probability that two draws have different colors?

The probability of drawing two balls of different colors for second draw would be P(1/6) if first draw results in two Green. The probability of drawing two balls of different colors for second draw would be P(4/15) if first draw results in two different colors.

What is probability of picking black marbles?

Step-by-step explanation: The probability of drawing a black marble from a bag is 1/4, and the probability of drawing a white marble from the same bag is 1/2.

What is the probability of choosing a yellow marble?

7/16
Event one: draw a yellow. Since there are 7 yellows out of a total of 16 marbles, the probability (desired outcomes over all possible outcomes) of drawing a yellow is 7/16.

What is the probability of choosing a red or green marble?

2 Answers By Expert Tutors The probability that the marble is either red or green is 0.8.

What is the probability of pulling out three different balls?

Without looking in the box, you pull out exactly three balls. What is the probability that the three balls will all be a different color (i.e. 1 red, 1 white and 1 blue)? The order of color does not matter, only that all three balls are different.

How to write the probability of pulling three different colors?

Using the laws of probability, we can see that the probability of each of the above events can be written as (changing the appropriate order, but we notice they are all the same) where, in the above expression, you must add the text “at first draw”, “at second draw”, “at third draw” respectively after B, W and R.

Is the probability of choosing two balls of the same color?

For the two of them to be in the same color I did 5/10 * 4/9 = 20/90 = 2/9 Is this the correct way and how would I calculate if they are of different colors because the 3 different colors are confusing me a bit. Thanks for contributing an answer to Mathematics Stack Exchange!

What’s the probability of drawing a blue ball?

Since they are disjoint, the probability that you are naming is which can be interpreted as the probability of drawing a blue ball or a white ball given than the first one was red. This is why now you cannot do what you state in point 3, since you have already considered the probability of getting a ball of blue or white color.